The following four wires are made of the same material Which of these will have the largest extension when the same tension is applied?
(a) length = 50 cm, diameter = 0.5 mm
(b) length = 100 cm, diameter = 1 mm
(c) length = 200 cm, diameter = 2 mm
(d) length = 300 cm, diameter = 3 mm.

Grade:12th pass

1 Answers

Aditi Chauhan
9 years ago

Ans (a) Y =T/1/?l/l??l = T * l/A*Y = T/Y * l/A

Here, T/Y is constant. Therefore, ?l ? l/A

l/A is largest in the first case.